6.111 Final Project Abstract

Chris Laas and Jacob Strauss

Monday, March 29, 1999

Our proposal for a 6.111 final project is an automated targeting system. It will consist of a motion detector module and a targeting module. The motion detector module will monitor a video stream and select regions of maximal change. The targeting module will choose a target and track it as long as it is visible. In addition, it will aim a laser pointer at the target by means of motors attached to rotating mirrors. The system will indicate when it is tracking a target and whether it has locked on to the target with its pointer. There will also exist a test function which will run the pointer through a sequence of motions preprogrammed into a PROM.